Calculate the size of the stellar aberration effect in Newtonian physics for
light coming straight down at speed c and a telescope moving straight
sideways at speed v. At what angle θ must we point the telescope to see
the star? Hint: In what direction is the light moving in the reference frame
of the telescope? Express θ in terms of v and c.
1
Expert's answer
2019-10-28T11:00:54-0400

Let's choose x-axis along the telescope speed "{\vec v}" and y-axis along the direction to the star (in accordance with the conditions of the problem angle between x-asix and y-axis is "\frac{\pi }{2}"). Then in the star's rest frame (we shall mark this values by apostrophe) the component of the velocity are "\vec u' = (0, - c)" . Then in the rest frame of the telescope the velocity of the light will be


"\vec u = \vec u' - \vec v = (-v, - c)"

Thus the tangent of the angle between the y-axis and the direction to the star is


"\tan (\frac{\pi }{2} - \theta ) = \frac{{{u_y}}}{{{u_x}}} = \frac{c}{v}"

Thus

"\tan (\theta ) = \frac{v}{c}"